OpenStudy (anonymous):

y varies inversely as twice x. When x = 2, y = 3. Find y when x = 6. 1 2 4 6

OpenStudy (anonymous):

y = k/(2x) Put the values of x and ya and solve for K (3) = k/(2(2)) 3 = k/4 k = 12 Write the equation again y = 12/(2x) Put x = 6 y = 12/(2(6)) y=12/12 y=1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Your welcome :)